OpenStudy (diamond17):

Need help on -2xt6≥18= Show work Please

OpenStudy (anonymous):

is that \[-2xt^6\ge18\] ?

OpenStudy (diamond17):

yes

OpenStudy (anonymous):

what do you need to solve for? x or t?

OpenStudy (diamond17):

i dont know this is what he ask for but i don't get it can u help please

OpenStudy (diamond17):

sorry the ? is this -2x+6≥18= i put the wrong thing

OpenStudy (anonymous):

Oh that's waay easier. \[-2x+6\ge 18\] you can add and subtract without worrying about the inequality so> \[-2x\ge 18-6\to -2x\ge 12\] then you divide by -2, requiring you to flip the inequality: \[x \le \frac{12}{-2}\to x\le -6\]

OpenStudy (diamond17):

Thanks

OpenStudy (diamond17):

can you help me graph it please

OpenStudy (diamond17):

freewilly922

OpenStudy (anonymous):

sorry, sure. on a number line go to the value -6 and cover with a line all the values LESS THAN OR EQUAL to -6 a dark dot over the line means including the point. the drawing is of a arrow to the left of -6 that has a filled dot at -6